OpenStudy (swagster23780):

The perimeter of a regular pentagon is represented by the expression 25c+15. What is the length of one side of the pentagon? Drag each expression into the table to show whether it is equivalent or not equivalent to the side length of the regular pentagon.

jhonyy9 (jhonyy9):

25c+15 is a perimeter of a penthagon - what mean 5 sides (25c+15)/5 = 5(5c+3)/5 = 5c+3 so this mean that 5c+3 will be the length of a side